### v1

# DONOHUE et al

Abstract
Longitudinal models that assume Gaussian residuals with proportional treatment group means provide direct estimates of the proportional treatment effect. However, we demonstrate that these models are biased and sensitive to the labeling of treatment groups. Typically, this bias favors the active group and inflates statistical power.
keywords:
proportional effect model; constrained longitudinal data analysis; cLDA
articletype: A Preprint
1 Background
Longitudinal models assuming Gaussian residuals with proportional treatment group means offer direct estimates of proportional treatment effect.1 In some instances, they seem to promise increased power2, and in other settings, they appear to be biased and relatively inefficient.3 The discrepant results are due to bias with the proportional effect model, which often favors the active group and artificially inflates power. The typical mixed model for repeated measures (MMRM) and the proportional MMRM differ only by how the treatment group difference is parameterized. To better understand the proportional treatment effect parameterization and its efficiency relative to the usual group mean difference, we can start by examining the cross-sectional setting.
2 Cross-sectional proportional effect model
Analogous to the longitudinal estimator defined in the tutorial, the cross-sectional proportional treatment effect estimator $\hat{\theta}$ is defined
$E(Y\mid x)=\hat{\beta}_{C}(1-\hat{\theta}x),$
where $x=0$ for the control group and $x=1$ for the active group. The mean in the control group is estimated by $\hat{\beta}_{C}$, and the active group estimate is $\hat{\beta}_{A}=\hat{\beta}_{C}(1-\hat{\theta})$, so that the estimated proportional reduction with active treatment is $\hat{\theta}=\frac{\hat{\beta}_{C}-\hat{\beta}_{A}}{\hat{\beta}_{C}}$. Expressing $\hat{\theta}$ this way shows its dependence on the control group mean $\hat{\beta}_{C}$. This implies that the model must assume that $\beta_{C}$ is sufficiently bounded away from zero.
Following similar notation, the familiar two-sample t-test estimator, $\hat{\delta}$, is defined
$E(Y\mid x)=\hat{\beta}_{C}+\hat{\delta}x.$
The additive treatment effect estimator is simply the difference $\hat{\delta}=\hat{\beta}_{A}-\hat{\beta}_{C}$, and in contrast $\hat{\delta}$ does not require that $\beta_{C}$ be bounded away from zero.
Figure 1: Simulated power (and type I error when $\delta$ = 0) for the proportional model and two-sample t-test based on 10,000 simulations. The control group mean is $\beta_{C}$ and $\delta$ is the mean group difference. With $\beta_{C}=0$, the proportional effect is not identifiable, and simulated power with the proportional test is about 5% regardless of $\delta$. With $\beta_{C}$ near zero ($\beta_{C}=-0.5$ or $-1$), the proportional test appears to have better power with $\delta>0$, and worse power with $\delta<0$. With $\beta_{C}$ far from zero ($\beta_{C}=-100$), power for the two methods is identical and the lines are overlapping. Residual variance was simulated to be one, sample size was 50 per group, and the proportional model was fit with R’s nonlinear least squares (nls) function.4
A simple simulation comparing statistical power for these cross-sectional models shows that the proportional model is sensitive to the direction of the effect (Figure 1). The proportional effect model appears to improve power, but only with $\delta>0$ and $\beta_{C}$ near zero. With $\delta<0$, the simple t-test generally dominates the proportional test. With $\beta_{C}=-100$ (bottom right) the power is identical with the two tests. With $\beta_{C}=0$ (top left), the treatment effect is not identifiable with the proportional model, and its power hovers around 5%. Figure 1 demonstrates that the proportional effect model is sensitive to the direction of the treatment effect and how groups are labeled.
Figure 2: Zipper plots 5 showing estimates and 95% confidence intervals for cross-sectional simulations with control group mean $\beta_{C}=-0.5$. The top row has a treatment effect of $\delta=0$, and demonstrates Type I error. The bottom row has a treatment effect of $\delta=0.3$, or equivalently, a proportional effect $\theta=0.6$. The intervals are sorted so that those associated with the largest standardized bias are toward the top of each panel, and only 25% of simulations with the largest bias are shown. While t-test estimates are symmetric about the true effect (left), the plots for the proportional test (right) reveal bias and asymmetric confidence interval widths. Vertical dashed lines are at the true value. Horizontal dashed lines are at the target rejection rate of 5%. Note that due to the numeric instability of the proportional model, p-values are often inconsistent with confidence interval coverage as can be seen with red intervals (p<0.05) which cover zero in the upper right panel.
Further inspection of the simulations shows the apparent improved power with the proportional test when $\beta_{C}$ is near zero is a result of bias. Figure 2 shows that with $\beta_{C}=-0.5$ and $\delta=0$, the proportional test tends to favor the active group (top right). Bias is also apparent with the proportional test and a treatment effect of $\delta=0.3$, or equivalently, a proportional effect $\theta=0.6$ (bottom right). We now explore the implications of this bias for longitudinal models.
3 Longitudinal proportional effect mode
Figure 3: Four longitudinal mean trend scenarios all with different proportional effects (ranging from 1/3 to 2) but equivalent linear effects (a difference between slopes of 0.5 outcome units per 18 months).
Figure 3 illustrates proportional effects ranging from 1/3 in Scenario A to 2 in Scenario D, but each panel has the same linear effect, a difference between slopes of 0.5 simulated outcome units per 18 months. All of these trends satisfy the proportional effect assumption. We simulate 10,000 trials under each of these scenarios with N=200 per group, residual variance 1.5, and random intercept variance 2. To each simulated dataset we fit (1) a constrained Longitudinal Data Analysis (cLDA) linear mixed-effect model assuming different slopes per treatment group and no group difference at baseline, and (2) a proportional effect cLDA (PcLDA) assuming a linear decline for the control group mean, and proportional treatment benefit. Because we are assuming a linear decline in the control group across all scenarios, both models correctly specify the temporal mean structure of the simulated data.
Scenario Model Power (%) Type I error (%) Proportion of rejections in favor of treatment under the null (%)
Scenario A Linear 88.3 4.9 47.4
Scenario A Proportional effect 91.5 4.5 71.5
Scenario B Linear 88.7 5.0 51.4
Scenario B Proportional effect 94.0 4.9 97.3
Scenario C Linear 88.2 5.0 49.4
Scenario C Proportional effect 94.8 5.2 100.0
Scenario D Linear 88.6 5.2 50.2
Scenario D Proportional effect 74.6 6.9 100.0
Table 1: Simulated power, Type I error, and proportion of rejections under the null that favor active treatment from 10,000 simulated trials under the mean trends shown in Figure 3.
Table 1 summarizes the longitudinal simulation results. The proportional effect model appears to have better power for Scenarios A, B, and C, but not Scenario D. Type I error is reasonably controlled except for the proportional effect model under Scenario D. But once again, under the null hypothesis, the proportional effect model is biased in favor of active treatment. Under Scenario A, when the proportional effect model rejects the null, it does so in favor of active treatment 72% of the time. This increases to 100% of the time for the other scenarios. The zipper plots5 in Figure 4 also demonstrate this bias in favor of treatment under the null hypothesis and tighter confidence intervals when estimates favor active treatment.
Figure 4: Zipper plots 5 showing estimates and 95% confidence intervals for longitduinal Type I error simulations. The intervals are sorted so that those associated with the smallest p-values are toward the top of each panel, and only the smallest 25% of p-values are shown. While linear model estimates are symmetric about the zero (left), the plots for the proportional effect model (right) reveal bias in favor of treatment with a disproportionate number of positive dots. Confidence intervals also appear to be narrower when proportional effect estimates favor treatment. Vertical dashed lines are at the true value, zero. Horizontal dashed lines are at the target rejection rate of 5%.
4 Discussion
These simulations demonstrate that the proportional effect parameterization is biased when the control group mean is near zero, and identical to the linear model when the control group mean is sufficiently far from zero. This bias occurs even when the proportional effect assumption is met. The apparent improved power of the longitudinal proportional effect model is likely due to bias in favor of active treatment. The proportional effect parameterization is sensitive to labeling of groups such that if we invert the group labels the model would be biased in favor of the control group. Biostatisticians should be aware of these fundamental issues with the proportional effect parameterization and avoid it whenever possible. Proportional effects and time savings 6 can always be examined based on linear models with the delta method (as mentioned in Wang et al. 20241), or the bootstrap, or by fitting with MCMC and interrogating the posterior distribution.
5 R Code for Cross-sectional Simulation
N <- 100
S <- 10000
BETA <- -c(seq(0, 1, by=0.5),100)
DELTA <- seq(-1, 1, by=0.1)
results <-
  parallel::mclapply(BETA, mc.cores = 4, function(Beta){
    set.seed(20240731)
    lapply(DELTA, function(Delta){
      lapply(1:S, function(sim){
        dd <- tibble(
          Tx = c(rep(0, N/2), rep(1,N/2)),
          Y = c(rnorm(N/2, Beta), rnorm(N/2, Beta+Delta)))
        tt <- lm(Y~Tx, data = dd)
        pt <- nls(Y ~ beta*(1-Tx*theta),
          data = dd,
          start = list(beta=coef(tt)[[1]], theta=0))
        ci.pt <- tryCatch(confint(pt), error = function(msg){
          return(tibble(‘2.5%‘=NA))})
        bind_rows(
          summary(tt)\$coef %>%
            as.data.frame() %>%
            rownames_to_column(’term’) %>%
            bind_cols(confint(tt)) %>%
            filter(term == ’Tx’) %>%
            mutate(Approach = ’t-test’) %>%
            rename(‘2.5%‘=‘2.5 %‘, ‘97.5%‘=‘97.5 %‘),
          summary(pt)\$coef %>%
            as.data.frame() %>%
            rownames_to_column(’term’) %>%
            bind_cols(ci.pt) %>%
            filter(term == ’theta’) %>%
            mutate(Approach = ’proportional test’)) %>%
          mutate(Simulation = sim, Beta = Beta, Delta = Delta)
      bind_rows()
      bind_rows()
      bind_rows()
save(results, file = ’results.rdata’)
6 R Code for Longitudinal Simulation
# define group-level means for four scenarios ----
long.means <- expand_grid(
  month = seq(0,18,by=3),
  tx = c(0,1),
  scenario = 1:4) %>%
  mutate(
    # control group means
    Y0 = case_when(
      scenario == 1 ~ -1.5/18 * month,
      scenario == 2 ~ -1.5/2/18 * month,
      scenario == 3 ~ -0.5/18 * month,
      scenario == 4 ~ -0.25/18 * month) %>%
        as.numeric(),
    # proportional treatment effects
    Y = case_when(
      scenario == 1 ~ Y0 * (1 - (1/3) * tx),
      scenario == 2 ~ Y0 * (1 - (2/3) * tx),
      scenario == 3 ~ Y0 * (1 - (1) * tx),
      scenario == 4 ~ Y0 * (1 - (2) * tx)) %>%
        as.numeric(),
    Tx = factor(tx, levels = 1:0, labels = c(’Active’, ’Control’)),
    Scenario = factor(scenario, levels = 1:4,
      labels = paste("Scenario", c("A", "B", "C", "D"))))
# expand group-level data to individual-level data ----
ind.long.means <- long.means %>%
  cross_join(tibble(id = 1:200)) %>%
  mutate(id = tx*200 + id)
N <- max(ind.long.means\$id)
# define proportional cLDA model ----
# covariates: month, tx (0 control, 1 active)
# parameters: Intercept, beta.month, theta
PcLDA <- function(month, tx, Intercept, beta.month, theta){
  Intercept + month*beta.month*(1-theta*tx)
# define function to fit models to simulated data ----
# this function is used for both power and type I error simulations
return.model.results <- function(data){
  lapply(1:4, function(scen){
    # fit linear model
    linear.model <- lme(Ysim ~ month + month:tx,
      random = ~ 1 | id, data = data %>filter(scenario == scen))
    # extract estimate and p-value for difference between slopes
    linear.result <- summary(linear.model)\$tTable %>%
      as.data.frame() %>%
      rownames_to_column(’term’) %>%
      bind_cols(intervals(linear.model)\$fixed) %>%
      filter(term == ’month:tx’) %>%
      mutate(Model = ’Linear’) %>%
      select(-Value)
    # use linear model fit for starting values for nonlinear model
    start1 <- fixef(linear.model)[c("(Intercept)", "month")]
    names(start1) <- c("Intercept", "beta.month")
    start1 <- c(start1, theta = 0)
    # fit non-linear proportional effect model
    nonlin.model <- nlme(Ysim ~ PcLDA(month, tx, Intercept,
      beta.month, theta) + b0,
      data = data %>filter(scenario == scen),
      fixed = Intercept + beta.month + theta ~ 1,
      random = list(b0 ~ 1),
      groups = ~ id,
      start = start1,
      control = nlmeControl(returnObject=TRUE))
    # extract proportional effect and p-value
    nonlin.result <- summary(nonlin.model)\$tTable %>%
      as.data.frame() %>%
      rownames_to_column(’term’) %>%
      bind_cols(intervals(nonlin.model)\$fixed) %>%
      filter(term == ’theta’) %>%
      mutate(Model = ’Proportional effect’) %>%
      select(-Value)
    # bundle and return results
    linear.result %>%
      bind_rows(nonlin.result) %>%
      mutate(scenario = scen)
  }) %>bind_rows()
# fix seeds for each of 10,000 simulated trials
set.seed(20241101)
SEEDS <- sample(1:2e9, size = 10000, replace = FALSE)
# run simulations ----
long.results <- parallel::mclapply(1:10000, mc.cores=10, function(sim){
  # set seed and simulate residuals and random intercepts
  set.seed(SEEDS[sim])
  dd <- tibble(
    id = 1:N,
    rand.int = rnorm(n=N, sd=2)) %>%
    right_join(ind.long.means, by = ’id’) %>%
    mutate(
      residual = rnorm(n=nrow(ind.long.means), sd=1.5),
      Y.power = Y + rand.int + residual,
      Y.null = Y0 + rand.int + residual
  # fit models
  power.results <- return.model.results(dd %>rename(Ysim = Y.power)) %>%
    mutate(Effect = ’Power’)
  null.results <- return.model.results(dd %>rename(Ysim = Y.null)) %>%
    mutate(Effect = ’Type I error’)
  bind_rows(power.results, null.results) %>%
    mutate(simulation = sim)
  bind_rows()
save(long.results, SEEDS, file = ’long.results.rdata’)